UT 27-Jan-2020 Engineering Half Night

Sunset 17:52

Twilight Ends: 19:12

Moon Sets 19:42

Night Center: 00:32 - Handover at 00:22 to accomodate instrument change
  • Eng Tasks
    • DONE OSCO - start of night - 30 min
      • run at leat 3x during the half-night (total time 1.5 hours)
      • Completed 2
    • AO/ESM on-sky checks (requires LUCI-1 operational - MOS beneficial to AO, not needed for ESM)
      • M67 - ESM on L1 / SL on L2 - visible 8:30-handover (R=12.2) - 0.75 hr
      • M67 - ESM on L1 / SL on L2 - visible 8:30-handover (R=16.3) - 0.75 hr
      • Arp 143 - ESM on L1 / SL on L2 - visible entire 1st half (R=14.7) - 0.75 hr
      • Collect data on individual stars
    • If clear - photometric standards for LUCI, MODS, LBCs - multiple airmasses needed (min 2)
      • G191B2b - visible entire first half - 0.5 hr
      • GSPC P247-U - visible entire first half - 0.5 hr
    • DONE Verify "Absorb Offset" warning works for LUCI (either or both) - IT 7965 - can verify with PCS GUI - 10min - can run on any script
      • I don't understand this IT - the behavior now is exactly as it has been for the last 3+years - its clear when the absorb has completed
    • ALERT! Speckle-Photometric tests (mixed L1-AO, L2-SL) - D. Thompson - 1hr - see note below
    • LBC/LUCI Mixed Mode test - full bino with dithers
      • Had to be run in pseudo-mono on UT 1/25/20, was working during restart: from IT 7500
      • The LUCI1 and LBCR scripts were generated from the OT, and the only change made after generating them was the add the line, <SYNCOFFSET>0,0,0,0,0 </SYNCOFFSET> to the syncoffset script (a 0 for each of the 5 dithers).
      • Could last TCS update have broken this?
      • Scripts: obsN:/home/lbto/Service_Observing/Eng_Dec_2019/Eng_Dec_MixedMode
  • D-time Targets
    • DONE TOO - A. Conrad/C. Veillet - 2hrs
    • QSOs (MODS) - O. Kuhn - 0.75hrs
  • Poor Weather Backups
    • Arp 187 LUCI-Bino or Mono - G200 zJ/HKspec 0.5"slit - 0.75 hour bino, 1.75 hr mono
    • Local Mergers CaT MODS - 0.75hr per target (3 targets)

OSA: J. Hill

ISp: G. Bechetti

Mtn Manager: R.Hansen

ISA: B. Rothberg, D. Thompson (Tucson)

AO: D. Miller, S. Walsh (Tucson)


Darks for INAF

UT 00:17 - D. Thompson started up the LUCIs - no issues

luciX.20200127.NNNN.fits

### darks_giannini_V900V960telluric_bino.xml
__LUCI1__ LUCI2 Comments________________
0001 0005 0001 0005 5x[5x12.0s] LIR/INT
0006 0010 0006 0010 5x[6x5.0s] LIR/INT
0011 0015 0011 0015 5x[2x30.0s] LIR/INT

### darks_rossiDDTspec+ima_bino.xml
__LUCI1__ LUCI2 Comments________________
0016 0020 0016 0020 5x[5x3.0s] LIR/INT
...skipped the 5x[6x5.0s] as it was done above
0021 0025 0021 0025 5x[1x200.0s] MER/NORM

# NOT DONE! Time to open the enclosure!
# ### darks_gargiuloVan0001+003_bino.xml
# LUCI1 LUCI2 Comments________________
# ...skipped the 5x[5x3.0s] as it was done above
# 0026 0030 0026 0030 5x[1x300.0s] MER/NORM

OSCO

UT 00:23 - setting up for opening up and OSCO

UT 00:31 - authorizing LUCI-1 only

Note: OT longslit scripts unlink if you change longslit to N30_Fieldstop...had to manually edit the script

UT 01:13 - pointing check done

UT 01:14- sending OSCO preset to HD 280675

UT 01:15 - LUCI-1 camera error - waiting for mask to go into FPU before mucking around with eng panels

-lots of +10 motions, commanded +691 steps, hit positive limit (now at 518) - making sure its there...going to N375 and N30

did not go to N375 new poition is -151 from 518

moving +690, now at +442 positive limit is on, camera init says N30

-looks like N30 - need to align it

--Next time use +50 steps to hit positive limit for N30

-error cannot align with pinholes

-doing it with box 3 in AFC tab (click X,Y click X,Y target)

UT 01:31 - struggling with collimation

-still waiting on collimation

UT 01:40 - sending OSCO preset HD 280675 L1#0026-0049

UT 01:55 - completed OSCO, ISAT appears to have shifted, team is re-calculating new position, in the meantime we will take another set of data

UT 01:57 - moving to OSCO target #2 HD 34066 L1#0050-

ISAT 901

UT 02:15 - J. Hill will send the track preset, i will center up manually taking images and then with the paddle

-slewing to target, first doing a pointing check

UT 02:21 - sending collimation script N375 no mask

UT 02:23 - LUCI-1 detector error - recovered with initialize then update the filters, camera and grating click update focus

L1 0070 -test ok

L1 0071 - alignng field stop

-BR. has used the paddle to center it up in LUCI

0073- verifying position to mathc rotator position

0074 - at 1025,1043 rotator center on LUCI i 1010, 1033

NOTE On telescope paddle: click OPTION to move CCD47 with paddle

-next step is to load up ACE Track script

-had to edit the LUCI script to have the AO Mag go from 30.0 to 12.0 (mag was 30.0 in the script) - any star without a mag is listed as 30.0 in the LUCI script (update that IT!)

--have to reset and make sure to click AO override on NSIGUI (2 takes to get this right)

0075 - acquisition N375 AO - 2.51x26sec +10 +10 offset

0076 -acquisition N375 AO - 2.51x6sec +0 +0 offset

0077 - acquisition N375 AO - 2.51x6sec +0 +0 offset N30 field stop in place "thru-slit" image

0078 - thru slit confirmation - looks good

0079- switching to N30 camera - fieldstop is off going to manually adjust it (pausing after exposure) - START OF SCIENCE

0080 - continuing with script - central few pixels are saturated in 2.51 seconds at K-band

0082- last of K-band without NCPA

0083 - NCPA is on - could be more saturated

0084 - mistake - loaded new script and forgot to correct 0.25" slit for N30....mask exchange is going on - loops have opened (2 mistakes for the price of one) - DO NOT USE THIS FRAME

UT 03:27 - putting n30 fieldstop back in

UT 03:34 -sending an offset seems to have recovered the ACE Track preset

UT 03:39 - recovered AO - closed loop - restarting at H-band

0085 - H-band first iamge - 0,0 offset but the target is near top of field stop - Image is OK (central pixels are also saturated)

1302,1687-- > need to offset -4.17 -9.94-4

Centered now

Going to Subframe "SW 1 895 1 256 2048"

UT 03:50 - offloading everything to the thin-shell --> losing AO loop

K-band is 45k counts in 1 frame - sub-framing to 128

Subframe "SW 1 959 1 128 2048"

UT 03:57 --we need to resend the preset, J. Hill will resend Track and I will do a re-acquisition with script (now cut down the exposure times since its bright)

UT 04:00 - may need to redo from start - > point and collimate to recover

DATA OBTAINED SO FAR:

5 K-band (1 NCPA), 1 H-band - full frame readout

ARGOS collimation script - if you pick a star or change the nominal configuration it doesn't write the guidestar into the script - so if I don't choose it from ARGOS-GS then it won't write it into the script

UT 04:16 - J. Hill sent track mode preset to field (NS) - taking manual luci image (N30 FS in turnout) H-band will paddle around to center star

0086 - initial acquisition of image - don't see it

moving -20 IE. -20 CA (i accidentally moved delta, but J. Hill corrected)

no sign of satellite

UT 04:31 - new satellite ephemeris generated - sending Track preset

UT 04:40 - taking data for PSF star - first image full frame - will center up on and then sub-frame

-elevation is lower than when data will be taken (airmass 1.778 vs 1.228)

UT 04:48 - this didn't work went to PA-40 should have been PA+40, plus target is setting, about to go back to Satellite (new ephemeris is generated)

Can't take a manual exposure on LUCI-1 restarting GEIRS on OBS4

-restarted readout manager on LMC

-Clicked "RESET" on readout manager - worked backlog read out (NOTE FOR FUTURE - this is the fix)

-fixed field stop

-We see it on CCD47

UT 05:11 - send ace track now

IGNORE frames 0086-0089

UT 05:18 - shell ripped

UT 05:22 - centered up in N30 fieldstop with N375 camera - tweaked IE CA, sending Ace Track NS

independent of AO setup - N375 image shows it well centerd, we are reconfiguring to N30 Camera and mask in FPU - can setup subframe to save time

-Field stop is offset, 300 pixels too south, FM4 motor 1 (Y) +1000 steps 9 steps per pixel (N30), +2000, +200, FM4 Motor 2 (X) 14steps per pixel (N30) should be -500 (X is negative to the right)

Telsvc -3" -15"

UT 05:39 - AO guys moving hexapod so our sub-frame didn't work - source moved out of X=128 sub-frame, but we are still more or less centered in the N30 fieldstop

UT 05:44 - edited script for N30 fieldstop in FPU, resending ACE Track for AO team (forgot to set it back to N30 Camera)

UT 05:48 - sent again, put back mask in turnout because we can't find target in CCD47 or LUCI

used paddles to center up for AO in CCD47, now putting mask in FPU and N30 camera in

-taking a manual exposure to setup sub-frame

UT 05:58 - centered up everything - subframe to 128x2048 - 0.15sec x 100

NOTE: -1" dither sends it off the frame 0095,0098 BAD (dithers off)

L1 0091 - 0105 K,H,J,z

Speckle-Photometry AO testing (D. Thompson)

UT 06:14 - clearing out NS setup, moving to pointing and collimation stars

-AO Ref star is 1/10th mm outside of the patrol field - OT shows the star is valid, AO cannot lock onto it

Other

UT 06:45 - LBCs powering up

UT 07:03 - Handing over early to AZ queue

-- BarryRothberg - 22 Jan 2020
Topic revision: r8 - 27 Jan 2020, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback